A sample of an ideal gas in a cylinder of volume 4.14 L at 298 K and 2.24 atm expands to 7.48 L by two different pathways. Path A is an isothermal, reversible expansion. Path B has two steps. In the first step, the gas is cooled at constant volume to 1.38 atm . In the second step, the gas is heated and allowed to expand against a constant external pressure of 1.38 atm until the final volume is 7.48 L . Calculate the work for path A. Calculate the work for path B.
Answer #1

Path A! Itt say idal gos = 1 moto W = nRT ln v2/11 49L = 1 mbl.x 8-314 584mb1-1 x 290 Klin este = 8987/35/3 = 14651587 212.95